OpenStudy (anonymous):

The sum of two consecutive odd numbers is 100. What is the smaller number?

OpenStudy (anonymous):

100/2 = 50 then add 1 to one of them and subtract one from the other. 49 + 51 = 100

OpenStudy (anonymous):

odd no. have a format of 2n+1 so two consecutive will be 2n+1, 2n+3 so just add them up 4n+4=100 find n!!

OpenStudy (anonymous):

that would make n=24 and that is not what it is